# RePrompt: Planning by Automatic Prompt Engineering for Large Language Models Agents

### v1

@@ Abstract
In this past year, large language models (LLMs) have had remarkable success in domains outside the traditional natural language processing, and people are starting to explore the usage of LLMs in more general and close to application domains like code generation, travel planning, and robot controls. Connecting these LLMs with great capacity and external tools, people are building the so-called LLM agents, which are supposed to help people do all kinds of work in everyday life. In all these domains, the prompt to the LLMs has been shown to make a big difference in what the LLM would generate and thus affect the performance of the LLM agents. Therefore, automatic prompt engineering has become an important question for many researchers and users of LLMs. In this paper, we propose a novel method, RePrompt, which does "gradient descent" to optimize the step-by-step instructions in the prompt of the LLM agents based on the chat history obtained from interactions with LLM agents. By optimizing the prompt, the LLM will learn how to plan in specific domains. We have used experiments in PDDL generation and travel planning to show that our method could generally improve the performance for different reasoning tasks when using the updated prompt as the initial prompt.
@@ 1 Introduction
Large language models (LLMs) have won significant success since the release of ChatGPT (OpenAI, 2022). In addition to traditional natural language tasks like summarization and sentiment analysis, LLMs have been shown to be effective in many domains that are closer to applications like code generation (Chen et al., 2023; Roziere et al., 2023), human-computer interaction (Li et al., 2023) and math problem solving (Wei et al., 2022; Yu et al., 2024). While pure LLMs are limited in their reasoning capability (Sun et al., 2023; Valmeekam et al., 2023; Chen et al., 2024), researchers have introduced tool-use to LLMs and built integrated systems, namely LLM agents, to enable the possibility of using LLM in even more general domains like robot controls (Wang et al., 2023a) and autonomous driving (Mao* et al., 2023).
Behind all of these successes, prompts are playing an important role. It has been shown that different prompts could lead to completely different success rates (Wei et al., 2022), and hence prompt engineering is often needed for each specific task. Because prompt engineering is difficult and time-consuming, automatic prompt engineering (APE) has emerged as a strategy, where the LLMs themselves write in the clear language they prefer (Zhou et al., 2023). With a limited number of trials, APE can efficiently converge to a robust prompt that performs better than a simple prompt on traditional natural language processing tasks (Zhou et al., 2023; Zhang et al., 2023).
However, most users still use primitive prompts or carefully hand-crafted prompts in their LLM. Besides the time and calculation needed for APE to converge to a good prompt, current processes of applying the APE-based framework have a few limitations that make it hard to outperform carefully hand-written prompts. Firstly, LLM does not know clearly what prompt works better for them. While LLM could write clear instruction sentences as the prompt, they do not show a better understanding of how different prompts will work on themselves when compared to humans. Additionally, because LLMs are fixed, they cannot learn the new paradigm people discovered, and when many people are trying on popular LLMs like GPT-4 (OpenAI, 2023a) or LLAMA-2 (Touvron et al., 2023), they get more ideas on how to write a good prompt. However, LLMs can only use their knowledge of how to prompt earlier models rather than how to prompt themselves. Furthermore, LLMs often cannot identify the correct solutions for a potential failure, and will only change to a slightly different prompt without improving. Specific to our objective in complex reasoning tasks, which are also known as LLM agents, many APE methods are primarily designed for straightforward question-answering (QA) tasks like HotPotQA (Yang et al., 2018) and fail to generalize to harder reasoning tasks.
@@
In this paper, we focus on a scenario where there is a specific reasoning task one wants to use LLMs for, with examples like people already choosing to use a specific OpenAI GPTs tool in ChatGPT (OpenAI, 2023b) to plan their travel or help in writing a code. In these domains, LLMs typically use a Chain-of-Thought (CoT) prompt with interactive procedures like ReAct (Yao et al., 2023b) and Reflexion (Shinn et al., 2023) to improve their performances. We propose a novel automatic prompt engineering method called RePrompt, which takes the common practices of using CoT and ReAct into consideration and uses the dialogue history from these results as the information for each prompt update. By summarizing the dialogue history and then analyzing how to improve the prompt sentence by sentence, we can successfully optimize the prompt, based on past history while not overfitting to corner cases. We further show how our proposed method could be seen as a standard fine-tuning procedure in ML where the changeable parameter (or input) is the initial prompt. An overview of our proposed framework is shown in Figure 1.
@@
We use experiments on Planning Domain Definition Language (PDDL) generation (Guan et al., 2023) and travel planning (Xie et al., 2024) to show that our methods can achieve a higher first-round success rate. These scenarios demonstrate the effectiveness of our approach when feedback is either expensive but accurate, or cheap but less accurate and informative.
@@
    Propose to use "gradient-based" prompt optimization in LLM agents.
@@
    Propose a summarization-based method to give specific instruction on how the current prompt could be further improved, and propose a novel guideline to optimize the prompt in a step-by-step format.
@@
    Using updated prompts, improve the results in multiple LLM agent benchmarks without fine-tuning LLM models on them.
@@ 2 Related Works
Our work lies in the intersection of prompt optimization and LLM for reasoning.
@@
In prompt optimization, many works proposed to optimize the prompt using differentiable tuning on soft prompts (Lester et al., 2021; Qin and Eisner, 2021), train auxiliary models as the optimizer (Hao et al., 2023; Deng et al., 2022; Zhou et al., 2023), or directly train the prompter themselves (Wang et al., 2023b). This line of work requires access to the model weights of the language models and is not generally applicable in the current era of using LLMs like GPT-4 (OpenAI, 2023a) and Claude-3 (Anthropic, 2024) through APIs. Another line of work chooses to use machine learning models to provide the approximated guidance on what prompt is better, either by using reinforcement learning (Shin et al., 2020; Zhang et al., 2023) or doing discrete manipulation with LLM feedbacks (Guo et al., 2023). There are also some other works in prompt optimization that propose a relatively general solution, like using beam search as the "gradient descent" optimizer (Pryzant et al., 2023; Tang et al., 2024). Our work is very close to the second group of work, and can be seen as a generalization from the current methods to domains that are more on reasoning domains.
@@
In LLM for reasoning, a key challenge identified by researchers is how to correctly use prompts to guide the LLM to generate useful auxiliary output that leads to a good final solution. Chain-of-Thought (CoT) (Wei et al., 2022) is the most commonly used prompt that can improve the performance of LLMs that consists of simply adding a fixed sentence of "Let’s think step by step.". Later on, Tree-of-thought (Yao et al., 2023a) and Graph of Thoughts (Besta et al., 2024) are also proposed as an extension from the simple line-based architecture of auxiliary output to tree and graph-structured output. Orthogonally, researchers have also found that utilizing the interaction capability of LLMs could also improve their performance. Yao et al. (2023b) proposed ReAct by letting the LLMs list some thoughts before proposing the actual actions. Reflexion (Shinn et al., 2023) prompts an LLM agent to reflect on the actions, and save the reflections in the memory to improve the efficiency. There are extensive amounts of works like self-refine (Madaan et al., 2023), RCI (Kim et al., 2023), and self-debugging (Chen et al., 2023) that use similar ideas of providing feedback as guidance in the reasoning-related task and strategically adapt the idea to fit the needs of specific domains. Our work considers this popular and useful workflow of iterating with feedback before finalizing the answer and optimizes the prompt based on the interaction history provided.
@@ 3 Methods
3.1 Preliminary on LLM agents
In this paper, we focus on LLM agents for reasoning tasks. Because LLM might not be correct in the first shot, in certain scenarios where feedback is still possible, prior approaches have proposed allowing a few interactions before the final answer is given by the LLM (Yao et al., 2023b; Shinn et al., 2023). In these cases, users provide the LLMs with some error information and let the LLMs try again with the additional information. This information does not necessarily provide any hint about the final solution but is an error message about why the current solution is not correct. For example, this information could be a typical Python runtime error message in code generation tasks. How to use specific prompts with error messages would mainly depend on different tasks. For example, in the widely used ReAct (Yao et al., 2023b), LLMs are required to provide thoughts on the current results before doing the next round, and these thoughts are often not a concrete error but may also include a short analysis that reaches a conclusion that the current action is good.
For LLM agents to focus on reasoning, the Chain-of-Thoughts (CoT) is one of the most popular methods used in LLMs. By adding a simple sentence of "Let’s do it step-by-step," LLMs will automatically consider outputting auxiliary steps before generating the final answer. Similar to how writing down the calculation procedure helps a high-school student solve math problems, the auxiliary steps are shown to be effectively helping the LLMs get a higher success rate in solving reasoning questions. From task to task, the auxiliary steps could differ significantly, and the exact steps needed in a task are called planning in LLM agents. A common practice is to let the LLM first generate these steps and then further choose whether each step should be executed by another LLM or external tools.
@@ 3.2 RePrompt
In this paper, we focus on addressing the planning part of LLM agents with prompt optimization. In this work, we focus on the problem of LLM agents where the task of the agents are known ahead. A typical example of this scenario is the different GPTs in OpenAI (OpenAI, 2023b).
As shown in Fig. 1, our method, RePrompt, is a prompt optimizer that is based on the interaction-based action generate process. Our method is quite similar to a typical machine learning training loop, but in our case, we are training the prompts to be fed into the model rather than the parameters of the LLMs. Given a specific small dataset of reasoning tasks used for training, we first let the LLMs generate their responses using the current prompt. This process needs to include some interaction schemes with some kind of feedback provider like ReAct or Reflexion , but we do not put any constraint on how this part should be done, and how accurate the feedback is. We call this process the act loop.
Then, we wait until a whole batch of this sort of chat history is collected, and we put the whole batch of chat history into an LLM model to summarize what is the most important focus point. The focus point could be a certain problem that appears quite often and leads to long iterations or specific suggestions (thoughts in the case of ReAct) that have shown to be effective in helping generate a good response. Typically, this information is already included in the chat history and does not need further analysis or summarization. We ask our summarization team to summarize different scenarios to remove scenarios’ specific information and suggestions, but also not summarize too much that will require further reasoning capability and could introduce hallucinations to the process if the LLM is not very good at the current task. Compared to previous work in automatic prompt engineering, including APE (Zhou et al., 2023) and ProTeGi (Pryzant et al., 2023), we do a summarization over a batch, and we prevent the prompt optimization from overfitting on a single outlier data point.
Then, with this summarized typical error, we use another LLM to update the actual prompt. We ask this optimizer LLM to follow the following rules while giving potential solutions:
@@
    The improvement should focus on the common prompt part rather than scenario-specific prompts that change from data point to data point. For example, in the task of formulation generation for PDDL, one of the most commonly appearing suggestions is to provide more information on what specific domain the LLM is trying to generate and more detailed background knowledge on what the domain is about. However, because we want to get a prompt that is general enough to solve all the PDDL formulation generation, such an update should not happen.
@@
    The improvement should prefer to identify whether the specific problem does occur in the given scenario. For example, suppose there is a certain budget one wants the solution provided by LLMs to satisfy, and the previous history shows that this budget constraint could be one of the main problems that lead to a failure in the solution. In that case, it should try to use the tools to calculate the cost for a typical plan. If it breaks the constraint, it could prioritize the budget constraint when getting a solution; otherwise, it should ignore this problem.
Based on the above rules, we ask the LLM to the following steps:
@@
    Choose the single solution that is the best. Unlike APE and some of the following work (Zhou et al., 2023; Deng et al., 2022), we do not ask the LLM to give a concrete number as the value of the sentence.
@@
    Output the final prompt that is a combination of the original prompt and the updated prompt.
After the steps, we will get an updated prompt, and we can continue to do more iterations, which can also be seen as more epochs as an analogy to training ML models, until the prompt has converged. This converged prompt will help improve the generated result in the first round. During test time, we only need to use the converged updated prompt and test it on the new test set. During the test, we do not require the exact same process to generate the response, e.g., the feedback generator can be removed completely from the Act procedure if it is quite expensive.
To help the optimizer address some common problems, we have currently manually encoded some of the common solutions in our optimizer prompt.
Note that in the optimization process, RePrompt only changes the step-by-step instruction phase rather than any other problem description or format requirement. This brings us to two possible formats of prompts that the algorithm will end up with:
@@
    If the current prompt is ReAct-like, which already includes a step-by-step instruction that gives a specific step, like Thought in ReAct, to include all the potential analysis, our prompt will converge to always update this thinking step by adding more and more hints on what to do to it. Our method becomes an algorithm that gives a more specific hint on what part the analysis should focus on compared to other prompt engineering work that introduces hints dynamically.
@@
    If the prompt is step-by-step, like solving a math or logical problem, then our algorithm is very likely to add more and more steps to the procedure generated by the planning. This will lead to a more concrete direction to try and focus on in the planning process, will guide the LLM in getting to the correct final answer, and will serve as the planner that breaks down the high-level task.
Note that, while we understand that in-context learning is very important for reasoning, we found it extremely challenging to update the examples to follow our instructions step-by-step perfectly all the time. And therefore, we choose not to change the examples at all. In most cases, the examples serve as a hint to the LLM on the output format and the related capabilities rather than a concrete guide on how to follow the step-by-step instructions, and we currently do not see any empirical drawback by not updating them.
For the very first training iterations, we are using the original prompt. While not every original prompt has detailed step-by-step instructions that our optimizer can build upon, we add an additional checker, which is not included in the workflow for simplicity, to convert the prompt into one with a step-by-step instruction. Specifically, we first use an LLM to check whether the current prompt already has step-by-step instructions, and if it does not, we manually add a sequence of instructions to the prompt that is fixed across domains to the location before any examples are given. The detailed sequence includes two steps: briefly analyze the problem and give the solution. This sequence is equivalent to a single CoT prompt (Wei et al., 2022), "let’s think step-by-step" in most reasoning tasks by introducing some necessary analysis while not providing any domain-specific human knowledge to the sequence.

### v2

@@ Abstract
In the past year, large language models (LLMs) have had remarkable success in domains outside the traditional natural language processing, and their capacity is further expanded into the so-called LLM agents when connected with external tools. In all domains, the prompt to the LLMs has been shown to make a big difference in what the LLM would generate and thus affect the performance of the LLM agents. Therefore, automatic prompt engineering (APE) has become an important question for many researchers and users of LLMs. However, previous works in APE rely on a final checker to evaluate the performance of the given prompt – a requirement that is hard to meet in the case of LLM agents, where intermediate feedback is easier to obtain, and the final evaluation could be expensive, inaccurate, or even missing. In this paper, we propose a novel method, RePrompt, which does a “gradient descent"-like approach to optimize the step-by-step instructions in the prompts given to LLM agents, based on the chat history obtained from interactions and reflections with LLM agents. By leveraging intermediate feedback, RePrompt can optimize the prompt without the need for a final solution checker. We evaluate our approach on PDDL generation, TravelPlanner, and Meeting Planning to show that our method could generally improve performance for different reasoning tasks.
@@ 1 Introduction
Large language models (LLMs) have won significant success since the release of ChatGPT (OpenAI, 2022). In addition to traditional natural language tasks like summarization and sentiment analysis, LLMs have been shown to be effective in many domains that are closer to applications like code generation (Chen et al., 2023; Roziere et al., 2023), human-computer interaction (Li et al., 2023) and math problem solving (Wei et al., 2022; Yu et al., 2024). While pure LLMs are limited in their reasoning capability (Sun et al., 2023; Valmeekam et al., 2023; Chen et al., 2024a), researchers have introduced tool-use to LLMs and built integrated systems, namely LLM agents, to enable the use of LLM in even more general domains (Wang et al., 2023a; Mao* et al., 2023; Xie et al., 2024).
Prompts play a crucial role in these successes, as variations in prompt design can lead to dramatically different success rates (Wei et al., 2022). Consequently, prompt engineering is often necessary for task-specific optimization. However, manual prompt engineering is both challenging and time intensive, motivating the development of automatic prompt engineering (APE), where LLMs generate prompts in their preferred natural language (Zhou et al., 2023). With some trials, APE can efficiently converge to a robust prompt, outperforming the original prompts that are often simple (Zhou et al., 2023; Zhang et al., 2023).
However, in complex LLM-agent tasks such as reasoning, APE is still under-studied, and most users still use primitive prompts or carefully hand-crafted prompts in their LLM. On the one hand, this is because LLM agents normally have high constraints on the output format, and previous APE methods can easily break the format requirement. On the other hand, existing APE methods rely highly on trying different prompts, checking the performance of each one, and finding the best one. However, there are numerous cases where an accurate final evaluator is highly costly, making it impractical for most people to use the evaluator frequently during training or, in our case, while optimizing the prompt. This is especially common in high-specialization domains that demand extensive domain knowledge, such as physics, chemistry, and other scientific disciplines. There are also scenarios where a final evaluation might be entirely absent, as seen in applications like ChatGPT, and particularly in GPT’s tools. In these situations, users interact with the application, sometimes providing intermediate feedback, but often leaving without clarifying whether they received a satisfactory response or abandoning the interaction due to dissatisfaction with the model’s ability to assist further. Without cheap prompt performance evaluators, existing APE methods cannot be applied to these scenarios.
@@
In this paper, we focus on a scenario where there is a specific reasoning task one wants to use LLMs for, while there is no ground-truth checker to check the correctness of the outputs. One example is the OpenAI GPTs tool in ChatGPT (OpenAI, 2023b) to plan their travel or help in writing a code. In these domains, LLMs typically use a Chain-of-Thought (CoT) prompt with interactive procedures like ReAct (Yao et al., 2023b) and Reflexion (Shinn et al., 2023) to improve their performance. We propose a novel automatic prompt engineering method called RePrompt, which takes the common practices of using CoT and ReAct into consideration and uses the dialogue history from these results as the information for each prompt update. By summarizing the dialogue history and then analyzing how to improve the prompt step-by-step, we optimize the prompt based on past history while not overfitting to corner cases. An overview of our proposed framework is shown in Figure 1.
@@
We benchmark our approach on Planning Domain Definition Language (PDDL) generation (Guan et al., 2023), TravelPlanner (Xie et al., 2024), and Meeting Planning Zheng et al. (2024) to show that our methods can achieve a higher first-round success rate, and our methods can be combined with different types of feedback generator.
@@
    Propose to use “gradient-based"-like prompt optimization in LLM agents.
@@
    Propose a summarization-based prompt optimization that focuses on optimizing steps in the prompt, and demonstrate that optimizing the steps is an efficient way of prompt optimization in LLM agents.
@@
    Our proposed method does not require a solution checker, and can be used in LLM-agents scenarios where such a checker is not available.
@@ 2 Related Works
Our work lies at the intersection of prompt optimization and LLM for reasoning.
@@
In prompt optimization, many works have proposed to optimize the prompt using differentiable tuning on soft prompts (Lester et al., 2021; Qin and Eisner, 2021), train auxiliary models as the optimizer (Hao et al., 2023; Deng et al., 2022; Zhou et al., 2023), or directly train the prompter themselves (Wang et al., 2023b). This line of work requires access to the model weights of the language models and is not generally applicable in the current era of using LLMs like GPT-4 (OpenAI, 2023a) and Claude-3 (Anthropic, 2024) through APIs. Another line of work chooses to use machine learning models to provide approximate guidance on which prompt is better, either by using reinforcement learning (Shin et al., 2020; Zhang et al., 2023; Chen et al., 2024b) or by discrete manipulation with LLM feedbacks (Guo et al., 2023). There are also some other works in prompt optimization that propose a relatively general solution, such as using beam search or Monte Carlo tree search as the "gradient descent" optimizer (Pryzant et al., 2023; Tang et al., 2024; Wang et al., 2024). Our work is very close to the second group of works, and can be seen as a generalization of the current methods to reasoning domains. Importantly, in our settings, we do not require ground-truth feedback that checks the general performance of the current prompt, which is more aligned with the applications of prompting for LLM agents in scenarios like GPTs OpenAI (2023b).
@@
In LLM for reasoning, a key challenge identified by researchers is how to correctly use prompts to guide the LLM to generate useful auxiliary output that leads to a good final solution. Chain-of-Thought (CoT) (Wei et al., 2022) is the most commonly used prompt that can improve the performance of LLMs that consists of simply adding a fixed sentence, such as "Let’s think step by step.". More recently, Tree-of-thought (Yao et al., 2023a) and Graph-of-Thoughts (Besta et al., 2024) were also proposed as an extension of the simple line-based architecture of auxiliary output to tree and graph-structured output. Orthogonally, researchers have also found that utilizing the interaction capability of LLMs could also improve their performance. Yao et al. (2023b) proposed ReAct by letting the LLMs list some thoughts before proposing the actual actions. Reflexion (Shinn et al., 2023) prompts an LLM agent to reflect on the actions, and save the reflections in the memory to improve efficiency. There are extensive other works like self-refine (Madaan et al., 2023), RCI (Kim et al., 2023), and self-debugging (Chen et al., 2023) that use similar ideas of providing feedback as guidance in the reasoning-related task and strategically adapt the idea to fit the needs of specific domains. Furthermore, recently, o1-like models Jaech et al. (2024) naturally integrate this self-correction process into the generation procedure. Our approach uses this iterative workflow that incorporates feedback before finalizing the answer, optimizing the prompt based on the provided interaction history. This process mitigates early-stage errors and, more importantly, reduces the inherent randomness seen in methods like ReAct and Reflexion, where certain scenarios may lack sufficient feedback. By summarizing recurring issues — such as budget balancing — our method, RePrmopt, ensures these concerns are systematically addressed within the prompt. This allows ReAct and Reflexion to focus on case-specific challenges, such as unexpected price surges for hotels in particular cities on specific days in the case of solving a TravelPlanner task.
@@ 3 Methods
3.1 Background on LLM agents
In this paper, we consider the problem of LLM agents for reasoning tasks. For LLM agents to focus on reasoning, the Chain-of-Thoughts (CoT) Wei et al. (2022) is one of the most popular methods used in LLMs. By adding a simple sentence of "Let’s do it step-by-step," LLMs will automatically outputting auxiliary steps before generating the final answer. Based on this, there has also been recent success in OpenAI o1 Jaech et al. (2024) and Deepseek-R1 et al. (2025) which further demonstrate that by introducing more steps before the answer, LLMs can solve harder problems.
Because LLM might not be correct in the first shot, prior approaches have proposed allowing a few interactions before the final answer is given by the LLM (Yao et al., 2023b; Shinn et al., 2023). In these cases, users provide the LLMs with some feedback and let the LLMs try again with the additional information. This information does not necessarily provide any hint about the final solution, but can be an error message about why the current solution is not correct. For example, this information could be a Python runtime error message in code generation tasks. How to use specific prompts with error messages would mainly depend on each specific task. For example, in the widely used ReAct (Yao et al., 2023b), LLMs are required to provide thoughts on the current results before doing the next round, and these thoughts are often not a concrete error but may also include a short analysis that reaches a conclusion that the current action is good. However, due to the difficulty of LLM agent’s tasks, checking how good a given prompt is could be expensive and unsuitable for regular in the prompt optimization process. As we mentioned earlier, this could be either very expensive in domains that require high intelligence like physics, or completely infeasible in applications like the web version of GPTs.
@@ 3.2 RePrompt
3.2.1 Overview
In LLM agents for reasoning tasks, we consider the task planning part of LLM agents with prompt optimization, where the tasks of the agents are known ahead of time, and the final solution checker is not available due to its cost.
As shown in Fig. 1, our method, RePrompt, is a prompt optimizer that is based on the interaction-based action-generating process. Our prompt optimization method is similar to a machine learning training loop, which iterates between getting the output based on the current parameters, calculating the loss based on the output, and optimizing the parameters based on the loss. But in our case, the parameters to be trained are the prompts to be fed into the LLM model, the model forward pass is replaced by the complete interaction-based action-generating process, which includes the feedback generator, and final components of RePrompt are the loss and optimizer, which are both LLMs instead of numerical calculation for the distance and the gradients.
Given a specific small dataset of reasoning tasks used for training, we first let the LLMs generate their responses using the current prompt. This process needs to include some interaction schemes with some kind of feedback generator like ReAct or Reflexion, but we do not put any constraint on how this part should be done, or how accurate the feedback is. We call this process the Actor.
We then wait until a complete batch of chat histories has been collected, at which point we input the entire batch into another large language model, which we call the summarizer, to summarize the primary focus point. This focus point might be a recurring issue that frequently prolongs iterations or specific suggestions (or “thoughts” in the case of ReAct) that have proven effective in producing high-quality responses. Typically, the essential information is already present in the chat history and does not require further analysis or summarization. Our summarizer is designed to capture key insights across various scenarios, omitting scenario-specific details and recommendations while avoiding overly broad summaries that would demand additional reasoning steps. Due to space constraints, we provide the prompt for this summarizer in Appendix D.
With the summarized typical errors identified, we leverage another LLM, the Prompt Optimizer, to refine the prompt accordingly. This optimizer LLM is instructed to adhere to the following rules:
@@ 3.2.1 Overview
    The refinement process should prioritize the common structural components of the prompt rather than scenario-specific elements that vary across data points. For instance, in the task of PDDL formulation generation, a frequent suggestion is to include additional details about the specific domain the LLM is targeting, along with more comprehensive background knowledge. However, since our goal is to construct a generalizable prompt capable of handling diverse PDDL formulation tasks, incorporating such domain-specific details would reduce its applicability and should therefore be avoided.
@@
    The improvement should prefer to identify whether the specific problem does occur in the given scenario. For example, suppose there is a certain budget one wants the solution provided by LLMs to satisfy, and the previous history shows that this budget constraint could be one of the main problems that lead to a wrong solution. In that case, the cost of a typical plan should first be approximated. If it breaks the constraint, it could prioritize the budget constraint when getting a solution; otherwise, it should ignore this problem.
Based on the above rules, we ask the LLM to do the following step-by-step, as detailed in Appendix D:
@@
    Choose the single solution that is the best. Unlike some of the existing work (Zhou et al., 2023; Deng et al., 2022), we do not ask the LLM to give a concrete number as the value of the prompt sentence.
@@
    Output the final prompt that combines the original prompt and the updated prompt.
To initialize the training process, we start with the original prompt. Although not all original prompts include detailed step-by-step instructions that our optimizer can leverage, we introduce an auxiliary checker (excluded from the main workflow for simplicity) to convert prompts into a structured, step-by-step format when needed. Specifically, we first use a language model to assess whether the current prompt already contains such instructions. If it does not, we manually append a standardized sequence of steps to the prompt, placing it just before any examples. This sequence comprises two primary steps: a brief problem analysis followed by the solution. This structure is functionally similar to a single "Chain of Thought" (CoT) prompt (Wei et al., 2022), “let’s think step-by-step” in most reasoning tasks by encouraging essential analysis without introducing domain-specific knowledge.
We integrate the specific rules and procedural steps mentioned above into our Optimizer prompt, detailed in the Appendix D due to the page limit. To assist the optimizer in handling common challenges, we have pre-encoded several frequently used solutions directly within the optimizer prompt. While these solutions could be discovered over multiple iterations, providing them upfront minimizes the number of iterations required for prompt refinement, accelerating the optimization process.
After the steps, we will get an updated prompt, and we can continue to do more iterations, which can also be seen as more epochs as an analogy to training ML models, until the prompt has converged and does not change with more iterations. This converged prompt will help improve the generated result in the first round, and also help assure common problems that can be fixed by the feedback generator to be resolved as early as possible. During test time, we only need to use the converged updated prompt and test it on the new test set. During the test, we do not require the exact same process to generate the response, e.g., the feedback generator can be removed completely from the Actor procedure if it is quite expensive.
Note that in the optimization process, RePrompt only changes the step-by-step instruction phase rather than any other problem description or format requirement specified in the prompt. This brings us to two possible formats of prompts that the algorithm will end up with:
@@
    If the current prompt is ReAct-like, or o1-like Jaech et al. (2024), which already includes a step-by-step instruction that gives a specific step, like Thought in ReAct, to include all the potential analysis, our prompt will converge to always update this thinking step by adding more and more hints on what to do to it. Our method becomes an algorithm that gives a more specific hint on what part the analysis should focus on compared to other prompt engineering work that introduces hints dynamically.
@@
    For step-by-step prompts, such as those involving mathematical or logical problem-solving, our algorithm progressively refines the generated procedure by incrementally adding steps. This iterative expansion enhances the concreteness of the planning process, providing a more structured pathway for the model to explore. By guiding the LLM toward the correct final answer, our approach effectively decomposes high-level tasks into granular substeps, improving both reasoning accuracy and interpretability.
While we understand that in-context learning is very important for reasoning, we found it extremely challenging to update the examples to follow our instructions step-by-step perfectly all the time. And therefore, we choose not to change the examples at all. In most cases, the examples serve as a hint to the LLM on the output format and the related capabilities rather than a concrete guide on how to follow the step-by-step instructions, and we currently do not see any empirical drawback by not updating them.